OpenStudy (anonymous):

Heres one I dont understand. (-2+SQRT-100)^2

OpenStudy (anonymous):

\[(-2+\sqrt{100})^2\]

OpenStudy (anonymous):

\[-2^2 + \sqrt{100}^2\] what's -2 * -2 and the sqrt of 100 squared just cancels the sqrt and the square thus making it 100

OpenStudy (anonymous):

4

OpenStudy (anonymous):

4+100i

OpenStudy (anonymous):

\[(-2+\sqrt{100})(-2+\sqrt{100})\] they don't cancel

OpenStudy (anonymous):

\[\sqrt{100} * \sqrt{100}\] doesn't cancle?

OpenStudy (anonymous):

just the sqrt

OpenStudy (anonymous):

100i?

OpenStudy (anonymous):

its -100

OpenStudy (anonymous):

\[(-2+\sqrt{100})(-2+\sqrt{100})\] i forgot you have to FOIL it out first outer inner last

OpenStudy (anonymous):

\[\sqrt{-100}\] i for got the negative

OpenStudy (anonymous):

yes so you can rewrite this as

OpenStudy (anonymous):

\[\sqrt{-1}\sqrt{100}\] where \[\sqrt{-1}=i\]

OpenStudy (anonymous):

you know that \[\sqrt{100}=10\]

OpenStudy (anonymous):

so now you have \[(-2+10i)(-2+10i)\]

OpenStudy (anonymous):

is it -100 or positive?

OpenStudy (anonymous):

negative sorry

OpenStudy (anonymous):

so foiling you et \[4-20i-20i+100i^2\] kowing that \[i^2=-1\] \[4-40i-100\] \[-96-40i\]

OpenStudy (anonymous):

i have the answer -96-40i

OpenStudy (anonymous):

that is the correct answer

OpenStudy (anonymous):

i thought you meant the middle i's cancel

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Thank you again for you helps!